OpenStudy (anonymous):

Use the fundamental trigonometric identities to simplify the expressions. Show all steps for full credit. HELP ME PLEASE 1.sec^2 x-tan^2 x= 2.cos^2 x(1+tan^2 x)= 3.secxcos(-x)-sin^2 x= 4.sec^2 x-1 / sin^2 x=

OpenStudy (anonymous):

1. \[\frac{ 1 }{\cos ^{2} x }\]-\[\frac{ \sin ^{2} x }{ \cos ^{2} x}\]=\[\frac{ (1-\sin ^{2}x) }{ \cos ^{2}x }\]=\[\frac{ \cos ^{2}x }{ \cos ^{2}x }\]=1 That's the general idea. just simplify the identities to sine and cosine

OpenStudy (anonymous):

thanks